### v2

# Detecting and protecting entanglement through nonlocality, variational entanglement witness, and nonlocal measurements

Abstract
We propose an innovative method to enhance the detection and protection of quantum entanglement, a cornerstone of quantum mechanics with applications in computing, communication, and beyond. While entanglement can be represented through nonlocal correlations detectable by the Clauser-Horne-Shimony-Holt (CHSH) inequality, this method does not fully capture all entangled states. To address this limitation, we introduce a variational entanglement witness (VEW) that optimizes the probabilities of detection and improves the efficiency of distinguishing between separable and entangled states. Additionally, we propose a novel nonlocal measurement framework that enables the assessment of both CHSH inequalities and the VEW while preserving the entanglement. Our approach enhances the reliability of entanglement detection while maintaining the entanglement of quantum states, offering significant advancements for quantum technologies.
I Introduction
Entanglement is a fundamental phenomenon where particles share correlated quantum states regardless of their spatial separation [1, 2]. It is crucial for many quantum technologies, such as quantum computing [3], quantum cryptography [4, 5], quantum communication [6, 7, 8], and quantum metrology [9, 10], among others. However, detecting and protecting entanglement from measurements is significant challenging due to the computationally intractable nature and the fragility of quantum states [11, 12].
One native approach is full quantum state tomography, which provides complete information about the quantum state [13, 14]. When the quantum state is reconstructed, one can evaluate the entanglement using criteria such as the Peres-Horodecki positive partial transpose (PPT) [15, 16] and concurrence [17]. However, tomography becomes impractical for large quantum systems due to its exponential scaling with the system size, making it highly resource-intensive. Recently, Elben et al. introduced a moments of the partially transposed density matrix (PT moments) protocol, using the first three PT moments to create a simple yet powerful bipartite entanglement test. The measurement was performed using local randomization, eliminating the need for full tomography [18].
An alternative method for detecting entanglement is using entanglement witnesses (EWs), which are operators that identify entanglement by measuring their expectation values [19, 20, 21, 22, 23, 24]. EWs offer a way to differentiate between entangled and non-entangled states without quantifying the degree of entanglement. The Bell theorem [25] and its associated inequalities, such as the Clauser-Horne-Shimony-Holt (CHSH) inequality [26, 27], fall within this category. These inequalities detect entanglement by revealing inconsistencies between quantum predictions and local realism. Violations of these inequalities provide evidence of entanglement, making them practical for distinguishing quantum states from classical ones. Although the CHSH inequality is widely used for this purpose [28, 29, 30, 31], it is crucial to note that entanglement can exist even when the inequality is not violated [32].
Recent advancements in machine learning have introduced promising methods for detecting quantum entanglement. Neural networks and support vector machines have proven effective in classifying quantum states as either entangled or separable [33, 34, 35, 23, 36, 37, 38, 39], detecting genuine multipartite entanglement [37], and developing entanglement witnesses [23, 20, 40, 39]. So far, convolutional neural networks have been particularly useful for analyzing entanglement patterns [41, 36]. These methods highlight the expanding role of machine learning in enhancing the detection and analysis of quantum entanglement.
However, these detection methods rely on multiple measurements of a quantum state and local measurements on spatially separable subsystems, which causes the collapse of the global wavefunction of the entire system. Therefore, there is a need for detection methods that can identify entanglement without destroying it.
In this paper, we introduce a variational entanglement witness (VEW) approach to detect entanglement when the CHSH inequality alone is not enough. While using these quantities for detecting entanglement is widespread, our approach provides an effective tool to confirm a quantum state entangled status. Optimizing VEW also improves the efficiency in distinguishing separable from entangled states. Moreover, we propose a nonlocal measurement framework to effectively measure the CHSH inequality and VEW, enabling both the detection and protection of entanglement.
Concretely, in a bipartite system shared with Alice and Bob, we first theoretically examine the entanglement via the violation of the CHSH inequality. We then apply a VEW to independently detect entanglement by training a parameterized witness operator. We extend the application of VEW to general pure and mixed states in two-dimensional systems and further to higher-dimensional systems. Finally, we propose a nonlocal measurement framework to measure the expectation values in the CHSH and VEW. This framework improves the reliability of entanglement detection and protects it from wave function collapse. We also use superconducting chips to simulate the nonlocal measurement and assess the post-measurement state to confirm the preservation of entanglement. This work not only advances our fundamental understanding of quantum entanglement but also supports quantum applications like secure communication and complex computations.
II Results
II.1 CHSH inequality and VEW
Suppose Alice and Bob share a bipartite system S represented by $|\psi\rangle$ as shown in Fig. 1(a). Each performs two experiments, with outcomes of either +1 or -1 on their respective parts. Alice measures operators $X$ and $Z$, while Bob measures $P$ and $Q$. The correlation between their measurements is given by the operator
$S_{\rm CHSH}=(X+Z)\otimes P+(X-Z)\otimes Q.$ (1)
In classical scenarios, $X,Z,P,Q\in\{\pm 1\}$ are random variables. In each run, $S_{\rm CHSH}$ can be either -2 or +2 according to the local hidden variable (LHV) theory. The expectation value follows the inequality
$|\langle S_{\rm CHSH}\rangle|\leq 2.$ (2)
In quantum mechanics, this inequality can be violated, ie., $|\langle S_{\rm CHSH}\rangle|>2$ [42, 43]. For example, let Alice and Bob share a quantum state
$|\psi\rangle=\cos\theta|00\rangle+e^{i\phi}\sin\theta|11\rangle,$ (3)
where the maximum entanglement occurs at $\theta=\pi/4$. To maximize the violation of the CHSH inequality, we choose the Pauli matrices for Alice as $X=|0\rangle\langle 1|+|1\rangle\langle 0|,Z=|0\rangle\langle 0|-|1\rangle\langle 1|$ and for Bob as $P=-(Z+X)/\sqrt{2},Q=(Z-X)/\sqrt{2}$ [43]. A direct calculation of Eq. (1) yields
$\langle S_{\rm CHSH}\rangle$ $=-\sqrt{2}\big{(}\langle ZZ\rangle+\langle XX\rangle\big{)}$
$=-\sqrt{2}\big{(}1+\cos\phi\sin 2\theta\big{)},$ (4)
where $ZZ$ abbreviates $Z\otimes Z$, and similarly for $XX$. See the detailed calculation in Appendix A. The CHSH inequality (2) is violated, ie., $|\langle S_{\rm CHSH}\rangle|>2$ at certain $\theta$ and $\phi$, as shown in Fig. 2(a). The aqua area marks the violation of the CHSH inequality, which also indicates the violation of the LHV model and demonstrates the nonlocality.
From this nonlocality, we can infer the entanglement in the quantum state $|\psi\rangle$. We confirm the entanglement using the PPT criterion and concurrence. Detailed calculations can be found in Appendix B. As depicted in Fig. 2(a), the maximum violation of inequality (2) corresponds to the highest level of entanglement, indicated by either minimum PPT or maximum concurrence. Thus, the violation of the CHSH inequality indicates the presence of entanglement.
While the nonlocal behavior can indicate entanglement, however, entanglement does not always imply nonlocality, meaning nonlocality cannot fully detect entanglement.
When the nonlocality fails to detect entanglement, an EW [44] is employed. An EW is a Hermitian operator $\mathcal{W}$ used to assess whether a quantum state $\rho$ is entangled. If $\text{Tr}(\mathcal{W}\rho)<0$, the state is identified as entangled; otherwise, if $\text{Tr}(\mathcal{W}\rho)\geq 0$, the state is considered non-entangled under this witness. In such a case, a different witness is needed for further evaluation.
An EW may require complete knowledge of the quantum state [1, 35]. For example, a general method to construct a witness involves the expression $\mathcal{W}=O-\min_{\rho\in\{\rho_{1}\otimes\rho_{2}\}}{\rm Tr}[O\rho],$ where $O$ is an arbitrary operator, or a witness based on the PPT criterion can be defined by $\mathcal{W}=(|v\rangle\langle v|)^{T_{A}}$, where $|v\rangle$ is the eigenvector of $\rho^{T_{A}}$ corresponding to the smallest eigenvalue, and $T_{A}$ denotes partial transposition [45]. For pure quantum states, $|\psi\rangle$, a projector witness for bipartite systems is given by $\mathcal{W}=\lambda_{\psi}^{2}\bm{I}-|\psi\rangle\langle|\psi|$, where $\lambda_{\psi}$ is the largest Schmidt coefficient [19]. In the case of two-qubit systems, a correlation-based EW can be expressed as $\mathcal{W}=\sum_{k}\sigma_{k}\otimes\sigma_{k},$ with $\sigma_{k}$ representing the Pauli matrices for $k\in\{0,x,y,z\}$ [46, 47].
Figure 1: (a) CHSH setup: an entangled bipartite state $|\psi\rangle$ is divided, with one subsystem sent to Alice and the other to Bob. Alice measures $X,Z$, while Bob measures $P,Q$. (b) A proposed nonlocal measurement model. In this model, two meters M1 and M2 are used to measure $ZZ$ and $XX$.
To find the most effective EW, here we propose a variational entanglement witness (VEW), which optimizes $\mathcal{W}$ to better detect entanglement for a given quantum state. The variational scheme proceeds as follows. First, the entanglement witness $\mathcal{W}$ is parameterized by $\bm{\upalpha}$, ie., $\mathcal{W}(\bm{\upalpha})$. The expectation value of this witness is then used as the cost function
$\mathcal{C}(\bm{\upalpha})=\text{Tr}\big{[}\mathcal{W}(\bm{\upalpha})\rho\big{]}.$ (5)
The optimization process is defined by
$\bm{\upalpha}^{*}$ $=\arg\min_{\{\bm{\upalpha}\}}\mathcal{C}(\bm{\upalpha})$ (6)
[…]
The objective is to find $\bm{\upalpha}^{*}$ that minimizes the cost function, ideally achieving a negative value for entangled states.
Figure 2: (a) Plot of $\langle S_{\rm CHSH}\rangle$ versus $\theta$ at different $\phi$ values. The CHSH inequality is violated in the aqua region where $|\langle S_{\rm CHSH}\rangle|>2$. Entanglement measures including the PPT criterion and concurrence are also shown for comparison. (b) Plot of the optimal entanglement witness $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle$.
In principle, VEW can be constructed from any operator, including Bell’s operator, however it requires perfect correlations, which are impractical experimentally. Here, we employ $S_{\rm CHSH}$ as the variational witness operator, thereby removing the requirement for prior knowledge of the quantum state and enabling direct measurement. The variational witness operator is defined by
$\mathcal{W}(\bm{\upalpha})=-\sqrt{2}\big{(}\alpha_{1}ZZ+\alpha_{2}XX\big{)},$ (8)
with the cost function is
$\mathcal{C}(\bm{\upalpha})=-\sqrt{2}\big{(}\alpha_{1}\langle ZZ\rangle+\alpha_{2}\langle XX\rangle\big{)}.$ (9)
The optimization process relies on gradient-free methods using the COBYLA optimizer. See detailed calculation in Appendix C.
II.1.1 Demonstration of VEW for a Bell state.
We first examine VEW for a Bell state given by $|\psi\rangle$ in Eq. (3). The results, illustrated in Fig. 2(b), effectively demonstrate the entanglement of the quantum state. At $\theta=0,\pi/2,$ and $\pi$, the expectation value $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle$ is zero, indicating zero in the PPT criterion and concurrence. At other points, $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle$ is negative, indicating the entanglement. Here, we achieve a 100% success rate, comparable to the performance of the machine learning approach [39]. For each point, we optimize VEW to obtain its minimum value. The minimum values are independent of one another, so the curves appear unsmooth. This observation holds for all $\phi$ except when $\phi=\pi/2$, because at this specific point, the expectation value $\langle XX\rangle=0$ and thus $\mathcal{C}(\bm{\upalpha})$ becomes a constant. According to the constraint in Eq. (7), this $\mathcal{C}(\bm{\upalpha})$ must vanish.
II.1.2 Demonstration of VEW for a general pure state.
For general cases, we examine a generic bipartite pure state represented by
$|\psi_{\rm gen}\rangle=a|00\rangle+b|01\rangle+c|10\rangle+d|11\rangle,$ (10)
where $a$, $b$, $c$, and $d$ are complex coefficients that satisfy the normalization condition $|a|^{2}+|b|^{2}+|c|^{2}+|d|^{2}=1$. The concurrence is calculated as
$C=2|ad-bc|.$ (11)
To evaluate numerically, we generate 100 random states, including 50 separable states ($C=0$) and 50 entangled states ($C>0$). VEW is then applied to classify the states as separable or entangled, with the results shown in Fig. 3. Refer to App. C for detailed information on the data generation process.
Figure 3(a) shows the set of random states, with blue dots representing separable states and red dots representing entangled states. The states are distributed randomly within an ellipse that illustrates the Hilbert space.
Figure 3(b) presents the classification results. The separable subspace is a convex subset that is nested within a larger convex set of all quantum states, while the entangled subspace is the region between them. Closed squares represent separable states (defined by $C=0$) and are placed in the separable subspace. Similarly, open squares represent entangled states and are assigned to the entangled subspace. Using VEW, these states are classified and assigned colors: blue means separable states with $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle\geq 0$ and red means entangled states with $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle<0$. The green dotted line indicates $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle=0$.
Correct classifications includes closed blue squares (separable states with positive VEW) and open red squares (entangled states with negative VEW). Misclassifications, on the other hand, appear as open blue squares (entangled states but positive VEW) and closed red squares (separable states but negative VEW). In this example of the random data set, VEW achieves 66% accuracy for separable states and 84% for entangled states. Moreover, states violating the CHSH inequality, marked by the smaller red dashed convex subset, are identified as entangled by both VEW and CHSH.
After all, we emphasize that both the CHSH inequality and VEW require nonlocal measurements of the expectation values $\langle ZZ\rangle$ and $\langle XX\rangle$. We will now present a framework for nonlocal measurements of $\langle ZZ\rangle$ and $\langle XX\rangle$ to confirm the CHSH inequality and VEW.
Figure 3: (a) A set of 100 random states, with 50 separable states shown as blue dots and 50 entangled states as red dots. The states are distributed within an ellipse representing the Hilbert space. (b) Classification results using VEW. The separable subspace is the blue ellipse, and the entangled subspace is the region between the red and blue ellipses. Closed squares indicate $C=0$ and opened squares are $C>0$. Blue colors indicate $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle\geq 0$ (separable), and red colors are $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle<0$ (entangled), separated by the green dotted line $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle=0$. The inset red dashed ellipse highlights the subspace where the CHSH inequality is violated.
II.2 Nonlocal measurement framework for measuring the CHSH inequality and VEW
To measure $X$ and $Z$ on the Alice’s side and $P$ and $Q$ on the Bob’s side, we need to measure the nonlocal products $Z\otimes Z$ and $X\otimes X$, or $ZZ$ and $XX$ for short. This is challenging because measuring noncommutative observables in a local state is impossible. To overcome this problem, we use entangled meters to couple to both Alice’s and Bob’s sides and readout the meters’ outcomes. The initial meters states are maximally entangled Bell states. Alice and Bob each couple their subsystem with the meters. After the interactions, they measure their meter in the $Z$ bases.
The nonlocal measurement model is shown in Fig. 1(b). To measure $\langle ZZ\rangle$ and $\langle XX\rangle$ simultaneously, we use two meters, $\rm M_{1}$ and $\rm M_{2}$. The meter states are given by Bell states as
$|\xi\rangle_{1}$ $=\dfrac{1}{\sqrt{2}}\Big{(}|\uparrow\uparrow\rangle+|\downarrow\downarrow\rangle\Big{)}$ (12)
$|\xi\rangle_{2}$ $=\dfrac{1}{\sqrt{2}}\Big{(}|\circ\!\times\rangle+|\times\!\circ\rangle\Big{)},$ (13)
[…]
Figure 4: (a) Quantum circuit for nonlocal measurement: System S consists of two qubits, $q_{0}$ and $q_{1}$, meter M1 consists of $q_{2}$ and $q_{3}$, and meter M2 consists of $q_{4}$ and $q_{5}$. (b) Probability $P(c_{0}c_{1}c_{2}c_{3})$ plotted for two cases: $(\theta,\phi)=(0,0)$ and $(\pi/4,0)$. (c) Plot of $\langle ZZ\rangle$ and $\langle XX\rangle$ versus $\theta$ at different $\phi$ values. The dotted curves are theoretical predictions, and the solid curves are from simulations. (d) Corresponding square error between simulations and theory plotted for $\langle XX\rangle$ case.
To measure $\langle ZZ\rangle$ (note that hereafter, we limit ourselves to the expectation values w.r.t the quantum state $|\psi\rangle$ in Eq. (3)), we apply the interaction $U_{1}$ between the system and M1 which are two CNOT (CX) gates as shown in Fig. 4(a). The measurement observables are represented by Kraus operators as
$M_{\mu}=\langle\mu|U_{1}|\xi\rangle_{1},\ \forall\mu\in\{\uparrow\uparrow,\uparrow\downarrow,\downarrow\uparrow,\downarrow\downarrow\},$ (14)
which gives
$M_{\uparrow\uparrow}$ $=M_{\downarrow\downarrow}=\dfrac{1}{\sqrt{2}}\big{(}|00\rangle\langle 00|+|11\rangle\langle 11|\big{)},$ (15)
$M_{\uparrow\downarrow}$ $=M_{\downarrow\uparrow}=0.$ (16)
Refer to Appendix D for detailed calculations. From these measurements, we obtain the probabilities $P_{\mu}=\langle\psi|M_{\mu}^{\dagger}M_{\mu}|\psi\rangle$ as
$P_{\uparrow\uparrow}=P_{\downarrow\downarrow}=\dfrac{1}{2};\ P_{\uparrow\downarrow}=P_{\downarrow\uparrow}=0.$ (17)
Then, the expectation value $\langle ZZ\rangle$ yields
$\langle ZZ\rangle=P_{\uparrow\uparrow}-P_{\uparrow\downarrow}-P_{\downarrow\uparrow}+P_{\downarrow\downarrow}=1.$ (18)
Even though $\langle ZZ\rangle=1$ regardless of $\theta$ and $\phi$ however, this value is not always known beforehand in other scenarios. Therefore, using $U_{1}$ is necessary to transfer information from the system to the meter $M_{1}$, enabling the measurement outcomes to reveal information about the system.
Similarly, we measure $\langle XX\rangle$ using meter M2, which interacts with the system through $U_{2}$, represented by two inverted CX gates as shown in Fig. 4(a). The measured operators are given by
$N_{\circ\circ}$ $=-N_{\times\times}=\dfrac{1}{2\sqrt{2}}\big{(}IX+XI\big{)},$ (19)
$-N_{\circ\times}$ $=N_{\times\circ}=\dfrac{1}{2\sqrt{2}}\big{(}IX-XI\big{)},$ (20)
and the corresponding probabilities are
$P_{\circ\circ}$ $=P_{\times\!\times}=\dfrac{1}{4}\big{(}1+\cos\phi\sin 2\theta\big{)},$ (21)
$P_{\circ\times}$ $=P_{\times\circ}=\dfrac{1}{4}\big{(}1-\cos\phi\sin 2\theta\big{)}.$ (22)
Finally, the expectation value $\langle XX\rangle$ is calculated as
$\langle XX\rangle=P_{\circ\circ}-P_{\circ\times}-P_{\times\circ}+P_{\times\times}=\cos\phi\sin 2\theta.$ (23)
Refer to detailed calculations in Append D. Consequently, both $\langle ZZ\rangle$ and $\langle XX\rangle$ can be measured through nonlocal measurements, thereby verifying the $S_{\text{CHSH}}$ inequality, and can also be used for optimizing the VEW.
II.3 Simulation on superconducting chip
We design a quantum circuit as shown in Fig. 4(a) for measuring $\langle ZZ\rangle$ and $\langle XX\rangle$. System S is $q_{0}q_{1}$, meter M1 is given by $q_{2}q_{3}$, and meter M2 is given by $q_{4}q_{5}$. The system state $|\psi\rangle$ is prepared by applying a quantum gate U3 onto $q_{0}$ and a CX gate onto $q_{0}q_{1}$, where
${\rm U}_{3}(\theta,\phi,\lambda)=\begin{pmatrix}\cos\frac{\theta}{2}&-e^{i\lambda}\sin\frac{\theta}{2}\\ e^{i\phi}\sin\frac{\theta}{2}&e^{i(\phi+\lambda)}\cos\frac{\theta}{2}\end{pmatrix},$ (24)
where we used $2\theta$ in Eq. (24) to get $|\psi\rangle$. Similarly, to prepare $|\xi\rangle_{1}$, we apply a Hadamard gate onto $q_{2}$ followed by a CX gate $q_{2}q_{3}$, and to prepare $|\xi\rangle_{2}$, we apply a Hadamard gate onto $q_{4}$, X gate onto $q_{5}$, followed by a CX gate $q_{4}q_{5}$. The interaction $U_{1}$ consists of two CX gates, while $U_{2}$ consists of two inverted CX gates as shown in the figure. Measure $q_{2},q_{3}$ in the $Z$ basis gives the outcome for $\langle ZZ\rangle$. To get $\langle XX\rangle$, we apply Hadamard gates onto $q_{4}$ and $q_{5}$, and measure on the $Z$ basis.
For the numerical experiment, we execute the quantum circuit using Qiskit simulation. For each data point, we run 10000 shots and obtain the classical probability $P(c_{0}c_{1}c_{2}c_{3})$, which is the outcome of the two meters M1 and M2, where $c_{0}c_{1}$ are the classical outcomes of M1 and $c_{2}c_{3}$ are the outcomes of M2.
In Fig. 4(b), we show $P(c_{0}c_{1}c_{2}c_{3})$ for several cases of $(\theta,\phi)$, including $(\theta,\phi)=(0,0)$ and $(\theta,\phi)=(\pi/4,0)$. First, we emphasize the bases denotation in Tab. 1 below.
Table 1: Bases denotation that used in the meters.
Bases
Meter 0 1
M1 $\uparrow$ $\downarrow$
M2 $\circ$ $\times$
[…]
$P_{ij}$ =$=\sum_{\{k,l\}\in\{\circ,\times\}}P(ijkl),$ (25)
$P_{kl}$ =$=\sum_{\{i,j\}\in\{\uparrow,\downarrow\}}P(ijkl),$ (26)
where $\{i,j\}\in\{\uparrow,\downarrow\}$ and $\{k,l\}\in\{\circ,\times\}$. Using these probabilities, we can calculate the nonlocal expectation values $\langle ZZ\rangle$ and $\langle XX\rangle$.
In Fig. 4(c), $\langle ZZ\rangle$ and $\langle XX\rangle$ are shown as functions of $\theta$ for different $\phi$ values. Simulation and theoretical results of these expectation values are compared and show good agreement. The corresponding square errors for $\langle XX\rangle$ are shown in Fig. 4(d) which are reasonable. These results show the effectiveness of nonlocal measurement in verifying the CHSH inequality and VEW.
II.4 Post-measurement quantum state
We derive the system state after these nonlocal measurements. The system (density) state after the first measurement gives [48]
$\rho_{1}=\sum_{\mu}M_{\mu}|\psi\rangle\langle\psi|M_{\mu}^{\dagger}=|\psi\rangle\langle\psi|,$ (27)
see Appendix E for detailed calculation. Next, we derive the system state after the second measurement. It gives
$\rho_{2}=\sum_{\nu}N_{\nu}\rho_{1}N_{\nu}^{\dagger}=\dfrac{1}{2}\begin{pmatrix}0&0&0&0\\ 0&1&\cos\phi\sin 2\theta&0\\ 0&\cos\phi\sin 2\theta&1&0\\ 0&0&0&0\end{pmatrix}.$ (28)
For example, Fig. 5(a) shows the tomography result of the final system state for $\theta=\phi=\pi/4$, which matches the theoretical calculation from Eq. (28).
Finally, we analyze the PPT criterion and concurrence of the final state to validate the protection of entanglement. For $\theta=k\pi/2$ or $\phi=\pi/2+k\pi$ for all $k\in\mathbb{N}$, the quantum state is $\rho_{2}={\rm diag}\big{(}0,1,1,0\big{)}/2$, where both the PPT criterion and concurrence are zero. This is depicted by the green dashed line in Fig. 5(b), ie., $\phi=\pi/2$. The polar plot of PPT criterion and concurrence against $\theta$ for different $\phi$ demonstrates that maximum entanglement occurs at $\theta=(2k+1)\pi/4$. These results indicate that entanglement is preserved under nonlocal measurement.
Figure 5: (a) Tomography of the final system state at $\theta=\phi=\pi/4$. (b) Polar plot of the PPT criterion and concurrence for the final state as functions of $\theta$ with different $\phi$ values.
II.5 Mixed state case
In this section, we examine $\langle S_{\rm CHSH}\rangle$ and its indication for the entanglement in mixed-state cases. We consider the Werner state as the system state
$\rho=p|\Psi^{-}\rangle\langle\Psi^{-}|+\dfrac{1-p}{4}\bm{I},$ (29)
where $0\leq p\leq 1$ and the Bell states (in the system bases) are defined by
$|\Phi^{\pm}\rangle$ $=\dfrac{1}{\sqrt{2}}\Bigl{(}|00\rangle\pm|11\rangle\Bigr{)},\text{and }$ (30)
$|\Psi^{\pm}\rangle$ $=\dfrac{1}{\sqrt{2}}\Bigl{(}|01\rangle\pm|10\rangle\Bigr{)}.$ (31)
Using the same meters M1 and M2 above, that means the same POVM as shown in Eqs. (D.7, D.8) and Eqs. (D.18, D.19). Then, we have
$P_{\uparrow\uparrow}$ $=P_{\downarrow\downarrow}={\rm tr}[\rho E_{\uparrow\uparrow}]=\dfrac{1-p}{4},$ (32)
$P_{\uparrow\downarrow}$ $=P_{\downarrow\uparrow}={\rm tr}[\rho E_{\uparrow\downarrow}]=\dfrac{1+p}{4},$ (33)
$P_{\circ\circ}$ $=P_{\times\times}={\rm tr}[\rho E_{\circ\circ}]=\dfrac{1-p}{4},$ (34)
$P_{\circ\times}$ $=P_{\times\circ}={\rm tr}[\rho E_{\circ\times}]=\dfrac{1+p}{4}.$ (35)
As a result, we have $\langle ZZ\rangle+\langle XX\rangle=-2p$, which implies $\langle S_{\rm CHSH}\rangle=2\sqrt{2}p$. The CHSH inequality (2) is violated when $1/\sqrt{2}<p\leq 1$. As previously stated, this violation region also exhibits entanglement.
Figure 6: (a) Plot of $\langle S_{\rm CHSH}\rangle$, PPT criterion, and concurrence as functions of $p$. (b) Plot of the optimal entanglement witness $\langle\mathcal{W}(\bm{\upalpha}^{*})\rangle$, and $\langle\mathcal{W}_{\rm Pauli}(\bm{\upalpha}^{*})\rangle$.
To further explore the entanglement behavior, we again calculate the PPT criterion and concurrence, with results shown in Fig. 6(a). First, in region I (sky blue area), $30\leq p\leq 1/$ 3, the system state is local and there is no entanglement, ie., the PPT criterion is positive and the concurrence is zero. In region II (orange area), $1/3<p\leq 1/\sqrt{2}$, the system is local but exhibits entanglement. Finally, in region III (light pink area), $1/\sqrt{2}<p\leq 1$, the system state is nonlocal and entangled. This analysis demonstrates that while nonlocal behavior can indicate entanglement, the reverse is not necessarily true. To further clarify, we provide a Venn diagram indicating the relationship between nonlocality and entanglement.
To fully detect entanglement, we use the VEW $\mathcal{W}(\bm{\alpha})$ as defined in Eq. (8) and compare its performance with the standard Pauli case, $\mathcal{W}_{\rm Pauli}(\bm{\alpha})=\alpha_{1}XX+\alpha_{2}YY+\alpha_{3}ZZ$. The results in Fig. 6(b) show that the Pauli-based witness often fails to distinguish between entanglement and separability. In contrast, the $\mathcal{W}(\bm{\alpha})$ case successfully detects entanglement for $p\geq 1/3$, ie., $\mathcal{W}(\bm{\alpha}^{*})<0$, and identifies separable states for $p<1/3$, ie., $\mathcal{W}(\bm{\alpha}^{*})\approx 0$. However, the differentiation between entangled and separable states becomes ambiguous around the critical value of $p=1/3$, making it challenging to conclusively determine the state.
Next, we derive the post-measurement states, which are given through
$\rho_{1}=\sum_{\mu}M_{\mu}\rho M_{\mu}^{\dagger}=\rho,$ (36)
and
$\rho_{2}=\sum_{\nu}N_{\nu}\rho_{1}N_{\nu}^{\dagger}=\begin{pmatrix}\frac{1+p}{4}&0&0&-\frac{p}{2}\\ 0&\frac{1-p}{4}&0&0\\ 0&0&\frac{1-p}{4}&0\\ -\frac{p}{2}&0&0&\frac{1+p}{4}\end{pmatrix}.$ (37)
See detailed calculation in Appendix F. Finally, we observe that the degree of entanglement in this state remains unchanged from the initial state $\rho$, indicating that entanglement persists under nonlocal measurement.
We next extend the use of VEW to mixed states beyond the Werner state by considering the Bell-diagonal state, a mixture of the four Bell states
$\rho=\sum_{i=1}^{4}\lambda_{i}|\Lambda_{i}\rangle\langle\Lambda_{i}|,$ (38)
where $\lambda_{i}$ are eigenvalues obeying $\sum\lambda_{i}=1$, and $|\Lambda_{i}\rangle$ are the Bell states defined in Eqs. (30, 31). The entanglement of the state is determined by the largest eigenvalue $\lambda_{i}$: the state is separable if $\mathrm{max}(\lambda_{i})\leq 0.5$, and entangled otherwise.
Similar to the pure state case in Fig. 3, we numerically generate 100 random states, evenly split between 50 separable and 50 entangled, by controlling ${\rm max}(\lambda_{i})$ as illustrated in Fig. 7(a). VEW is then applied for classification, with the results shown in Fig. 7(b). VEW achieves classification accuracies of 66% for separable states and 70% for entangled states. In this example dataset, none of the entangled states violate the CHSH inequality. See also Fig. 9(b).
II.6 Demonstration of VEW for high-dimension systems
We have discussed various applications of the proposed VEW for two-dimensional bipartite systems with pure and mixed states. Now, we apply it to high-dimensional bipartite systems.
Figure 7: Classification of mixed states using VEW, similar to the pure states case in Fig. 3. (a) States are generated with ${\rm max}(\lambda_{i})\leq 0.5$ for separable states and ${\rm max}(\lambda_{i})>0.5$ for entangled states. (b) Classification results using VEW.
We consider a bipartite system of two qudit subsystems in the Hilbert space $H=\mathbb{C}_{A}^{d}\otimes\mathbb{C}_{B}^{d}$, with canonical bases $\{|i,i\rangle\equiv|i_{A}\rangle\otimes|i_{B}\rangle\}_{i=0}^{d-1}$, where $d$ is the dimension of each qudit. The Bell inequality generalizes to the Collins-Gisin-Linden-Massar-Popescu (CGLMP) inequalities for higher-dimensional systems [49]. We define the CGLMP operator $S_{\rm CGLMP}$ as
$S_{\rm CGLMP}=(X_{d}+Z_{d})\otimes P_{d}+(X_{d}-Z_{d})\otimes Q_{d},$ (39)
where the shift and phase operators $X_{d}$ and $Z_{d}$ are
$X_{kl}=\begin{cases}1&\text{if }l=(k+1)\mod d,\\ 0&\text{otherwise},\end{cases}$ (40)
$Z_{kl}=\begin{cases}\exp\left(\frac{2\pi i}{d}\cdot k\right)&\text{if }k=l,\\ 0&\text{otherwise},\end{cases}$ (41)
and $P_{d}=-(Z_{d}+X_{d})/\sqrt{2},Q_{d}=(Z_{d}-X_{d})/\sqrt{2}$.
The common used EW $\mathcal{W}_{d}$ is derived from $S_{\rm CGLMP}$ as
$\mathcal{W}_{d}=2\textbf{I}_{d}-S_{\rm CGLMP},$ (42)
where $\textbf{I}_{d}$ is the identity operator in $d$-dimensional space. We also define the VEW as
$\mathcal{W}_{d}(\bm{\upalpha})=-\sqrt{2}\left(\upalpha_{1}Z_{d}\otimes Z_{d}+\upalpha_{2}X_{d}\otimes X_{d}\right),$ (43)
and the cost function $\mathcal{C}_{d}(\bm{\upalpha})=\langle\mathcal{W}_{d}(\bm{\upalpha})\rangle$. The optimization problem is given by
$\bm{\upalpha}^{*}$ $=\arg\min_{\{\bm{\upalpha}\}}\mathcal{C}(\bm{\upalpha})$ (44)
[…]
Figure 8 shows the numerical results for several quantum states, including the separable state $|\psi\rangle=|0,0\rangle$ and two entangled states: $|\psi\rangle=\sum_{i}|i,i\rangle/\sqrt{d}$ (maximum entangled state) and $|\psi\rangle=(|0,0\rangle+|d-1,d-1\rangle)/\sqrt{2}$ (partial entangled state). We compare the performance of EW $\mathcal{W}_{d}$ (solid curves) and VEW $\mathcal{W}_{d}(\bm{\upalpha}^{*})$ (dotted curves). Although the EW does not distinguish between entangled and separable states (all $\mathcal{W}_{d}$ are negative), the VEW successfully classifies them ($\mathcal{W}_{d}(\bm{\upalpha}^{*})=0$ for separable state and $\mathcal{W}_{d}(\bm{\upalpha}^{*})<0$ for entangled states), demonstrating its advantage in detecting entanglement in high-dimensional systems.
Figure 8: Classification of quantum states in high-dimensional systems using EW $\mathcal{W}_{d}$ (solid curves) and VEW $\mathcal{W}_{d}(\bm{\upalpha}^{*})$ (dotted curves): EW fails to distinguish between entangled and separable states, whereas VEW successfully classifies them.
III Conclusion
We made significant progress in detecting and protecting quantum entanglement using nonlocality, variational entanglement witness (VEW), and nonlocal measurements. While traditional methods like violations of the CHSH inequality are effective, they do not cover all scenarios of entanglement detection. By introducing VEW, we addressed the limitations of these traditional methods, and offered a more comprehensive approach to identifying entanglement. We also proposed a nonlocal measurement framework for measuring the CHSH operator and VEW. Our findings emphasize the crucial role of nonlocal measurements in detecting and maintaining entanglement, which are essential for the functionality of quantum technologies. This work not only advances the understanding of quantum entanglement but also contributes to the practical development of more robust quantum computing, communication, and sensing systems.
Acknowledgements.
This paper is supported by JSPS KAKENHI Grant Number 23K13025. The code is available at: https://github.com/echkon/nonlocalMeasurement
Appendix A CHSH inequality
In this Appendix, we derive details of the CHSH inequality. We start from Eq. (1) in the main text as
$S_{\rm CHSH}=(X+Z)\otimes P+(X-Z)\otimes Q.$ (A.1)
Using $P=-(Z+X)/\sqrt{2},\ Q=(Z-X)/\sqrt{2}$, we have
$S_{\rm CHSH}$ $=-(X+Z)\otimes\dfrac{(Z+X)}{\sqrt{2}}+(X-Z)\otimes\dfrac{(Z-X)}{\sqrt{2}}$
$=-\sqrt{2}(XX+ZZ).$ (A.2)
Now, the expectation value gives
$\langle S_{\rm CHSH}\rangle$ $=-\sqrt{2}\langle\psi|\big{(}XX+ZZ\big{)}|\psi\rangle$
$=-\sqrt{2}\big{(}\langle 00|\cos\theta+e^{-i\phi}\sin\theta\langle 11|\big{)}$
[…]
×$\hskip 28.45274pt\times\big{(}\cos\theta|00\rangle+e^{i\phi}\sin\theta|11\rangle\big{)}$
$=-\sqrt{2}(\cos^{2}\theta+\sin^{2}\theta+\dfrac{1}{2}(2\cos\phi)\sin 2\theta)$
$=-\sqrt{2}(1+\cos\phi\sin 2\theta)$ (A.3)
Appendix B Entanglement measures
We discuss various entanglement measures here, including the PPT criterion [15, 16] and concurrence [17] for bipartite systems. Let $\rho$ be a density matrix of an arbitrary mixed state of the two-qubit system AB. Its partial transpose (with respect to the B party) is defined as
$\rho^{T_{B}}=(I\otimes T)[\rho].$ (B.1)
If $\rho^{T_{B}}$ has a negative eigenvalue, $\rho$ is guaranteed to be entangled.
For concurrent, we first derive $\rho_{C}$ as
$\rho_{C}=(Y\otimes Y)\rho^{*}(Y\otimes Y),$ (B.2)
where $\rho^{*}$ is the complex conjugate of $\rho$. Let $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}$ are eigenvalues of $\rho_{C}$, the concurrence is defined by
$\mathcal{C}(\rho)=\max\big{\{}0,\lambda_{1}-\lambda_{2}-\lambda_{3}-\lambda_{4}\big{\}}.$ (B.3)
The bipartite system is entangled if $\mathcal{C}(\rho)>0$ and the maximum of $\mathcal{C}(\rho)$ means the maximum of entanglement.
Appendix C Variation entanglement witness
In this section, we outline the optimization process for the variational entanglement witness. The cost function is defined as
$\mathcal{C}(\bm{\upalpha})=\text{Tr}\big{[}\mathcal{W}(\bm{\upalpha})\rho\big{]},$ (C.1)
where $\mathcal{W}(\bm{\upalpha})$ is a Hermitian operator parameterized by $\bm{\upalpha}$, and $\rho$ is the quantum state of interest.
To minimize the cost function, we use the gradient-free COBYLA optimizer. This optimization yields the final parameters $\bm{\upalpha}^{*}$, which minimize the cost function $\mathcal{C}(\bm{\upalpha})$ and represent the optimal parameters for minimizing $\text{Tr}[\mathcal{W}(\bm{\upalpha})\rho]$.
For pure states, where $\mathcal{W}(\bm{\upalpha})=-\sqrt{2}(\alpha_{1}ZZ+\alpha_{2}XX)$ and $\rho=|\psi\rangle\langle\psi|$, we use the set of separable states $\{|00\rangle\langle 00|,|01\rangle\langle 01|,|10\rangle\langle 10|,|11\rangle\langle 11|\}$. The initial parameters are $\bm{\upalpha}=[4.0,0.0]$.
For mixed states, we use both the witness operator $\mathcal{W}(\bm{\upalpha})$ and a Pauli-based witness operator $\mathcal{W}_{\rm Pauli}(\bm{\upalpha})=\alpha_{1}XX+\alpha_{2}YY+\alpha_{3}ZZ$. The quantum state $\rho$ is the Werner state, as defined in Eq. (29). In addition to the previous separable states, we include $\rho(p=0)=\bm{I}/4$. The initial parameters for the Pauli-based case are chosen randomly.
Additional data for general random pure and mixed states. For general pure states, 100 random states were generated according to Eq. (10), consisting of 50 separable states ($C=0$) and 50 entangled states ($C>0$). The PPT, concurrence, and $\langle S_{\rm CHSH}\rangle$ values were analyzed, as shown in Fig. 9(a). Similarly, Fig. 9(b) presents results for 100 random mixed states, generated from Eq. (38), with 50 states satisfying $\text{max}(\lambda_{i})\leq 0.5$ and 50 states satisfying $\text{max}(\lambda_{i})>0.5$. These data are used to examine VEW, as detailed in Figs. 3 and 7 of the main text.
Figure 9: (a) Generating of 100 random pure states, with 50 states having $C=0$ and 50 states having $C>0$. For each state, the PPT, concurrence, and $\langle S_{\rm CHSH}\rangle$ values are displayed. (b) A similar process was applied to mixed states, with 50 states having ${\rm max}(\lambda_{i})\leq 0.5$ and 50 states having ${\rm max}(\lambda_{i})>0.5$ These data are used to examine VEW, as shown in Figs. (3, 7) of the main text.
Appendix D Nonlocal measurement
In this section, we derive detailed calculations of nonlocal measurement of $\langle ZZ\rangle$ and $\langle XX\rangle$.
D.1 Nonlocal measurement of $\langle ZZ\rangle$
The measurement of $\langle ZZ\rangle$ is given through meter M1 is initially prepared in $|\xi\rangle_{1}=\dfrac{1}{\sqrt{2}}\big{(}|\uparrow\uparrow\rangle+|\downarrow\downarrow\rangle\big{)}$. First, the interaction between the system and the meter M1 is given through $U_{1}$, which are CX gates as
${\rm CX}_{q_{0}q_{2}}\ {\rm CX}_{q_{1}q_{3}}$ $=\Big{(}|0\rangle\langle 0|_{q_{0}}\otimes I_{q_{2}}+|1\rangle\langle 1|_{q_{0}}\otimes X_{q_{2}}\Big{)}\Big{(}|0\rangle\langle 0|_{q_{1}}\otimes I_{q_{3}}+|1\rangle\langle 1|_{q_{1}}\otimes X_{q_{3}}\Big{)}$
$=\Big{(}|00\rangle\langle 00|_{q_{0}q_{1}}\otimes I_{q_{2}}I_{q_{3}}+|01\rangle\langle 01|_{q_{0}q_{1}}\otimes I_{q_{2}}X_{q_{3}}+|10\rangle\langle 10|_{q_{0}q_{1}}\otimes X_{q_{2}}I_{q_{3}}+|11\rangle\langle 11|_{q_{0}q_{1}}\otimes X_{q_{2}}X_{q_{3}}\Big{)},$ (D.1)
[…]
$U_{1}|\xi\rangle_{1}$ $=\Big{(}|00\rangle\langle 00|\otimes II+|01\rangle\langle 01|\otimes IX+|10\rangle\langle 10|\otimes XI+|11\rangle\langle 11|\otimes XX\Big{)}\otimes\dfrac{1}{\sqrt{2}}\big{(}|\uparrow\uparrow\rangle+|\downarrow\downarrow\rangle\big{)}\big{)}$
$=\dfrac{1}{\sqrt{2}}\Big{(}|00\rangle\langle 00|+|11\rangle\langle 11|\Big{)}\otimes\Big{(}|\uparrow\uparrow\rangle+|\downarrow\downarrow\rangle\Big{)}+\Big{(}|01\rangle\langle 01|+|10\rangle\langle 10|\Big{)}\otimes\Big{(}|\uparrow\downarrow\rangle+|\downarrow\uparrow\rangle\Big{)}.$ (D.2)
The measure observables are given in Kraus operators as $M_{\mu}=\langle\mu|U_{1}|\xi\rangle_{1},\forall\mu\in\{\uparrow\uparrow,\uparrow\downarrow,\downarrow\uparrow,\downarrow\downarrow\}$, where
$M_{\uparrow\uparrow}$ $=\langle\uparrow\uparrow|U_{1}|\xi\rangle_{1}=\dfrac{1}{\sqrt{2}}\big{(}|00\rangle\langle 00|+|11\rangle\langle 11|\big{)},$ (D.3)
$M_{\uparrow\downarrow}$ $=\langle\uparrow\downarrow|U_{1}|\xi\rangle_{1}=\dfrac{1}{\sqrt{2}}\big{(}|01\rangle\langle 01|+|10\rangle\langle 10|\big{)},$ (D.4)
$M_{\downarrow\uparrow}$ $=\langle\downarrow\uparrow|U_{1}|\xi\rangle_{1}=\dfrac{1}{\sqrt{2}}\big{(}|01\rangle\langle 01|+|10\rangle\langle 10|\big{)},$ (D.5)
$M_{\downarrow\downarrow}$ $=\langle\downarrow\downarrow|U_{1}|\xi\rangle_{1}=\dfrac{1}{\sqrt{2}}\big{(}|00\rangle\langle 00|+|11\rangle\langle 11|\big{)}.$ (D.6)
We next calculate the positive operator-valued measure (POVM) $E_{\mu}=M_{\mu}^{\dagger}M_{\mu}$, which give
$E_{\uparrow\uparrow}=E_{\downarrow\downarrow}=M_{\uparrow\uparrow}^{\dagger}M_{\uparrow\uparrow}=\dfrac{1}{2}\big{(}|00\rangle\langle 00|+|11\rangle\langle 11|\big{)},$ (D.7)
$E_{\uparrow\downarrow}=E_{\downarrow\uparrow}=M_{\uparrow\downarrow}^{\dagger}M_{\uparrow\downarrow}=\dfrac{1}{2}\big{(}|01\rangle\langle 01|+|10\rangle\langle 10|\big{)}.$ (D.8)
And the probabilities yield
$P_{\uparrow\uparrow}$ $=\langle\psi|E_{\uparrow\uparrow}|\psi\rangle$
$=\big{(}\cos\theta\langle 00|+e^{-i\phi}\langle 11|\big{)}\dfrac{1}{2}\big{(}|00\rangle\langle 00|+|11\rangle\langle 11|\big{)}$
×$\hskip 85.35826pt\times\big{(}\cos\theta|00\rangle+e^{i\phi}|11\rangle\big{)}$
$=\dfrac{1}{2},$ (D.9)
and
$P_{\uparrow\downarrow}$ $=P_{\downarrow\uparrow}=\langle\psi|E_{\uparrow\downarrow}|\psi\rangle=0.$ (D.10)
$P_{\downarrow\downarrow}$ $=\langle\psi|E_{\downarrow\downarrow}|\psi\rangle=\dfrac{1}{2}.$ (D.11)
Finally, we obtain the expectation value
$\langle ZZ\rangle=P_{\uparrow\uparrow}-P_{\uparrow\downarrow}-P_{\downarrow\uparrow}+P_{\downarrow\downarrow}=1,$ (D.12)
as shown in Eq. (18) in the main text.
D.2 Nonlocal measurement of $\langle XX\rangle$
Similar, to measure $\langle XX\rangle$, we use meter M2 with the initial state $|\xi\rangle_{2}=\dfrac{1}{\sqrt{2}}\big{(}|\circ\!\times\rangle+|\times\!\circ\rangle\big{)}$. The interaction $U_{2}$ is the inverted CX gates
$\bar{\rm CX}_{q_{0}q_{4}}\bar{\rm CX}_{q_{1}q_{5}}$ $=\big{(}I_{q_{0}}\otimes|\circ\rangle\langle\circ|_{q_{4}}+X_{q_{0}}\otimes|\times\rangle\langle\times|_{q_{4}}\big{)}\big{(}I_{q_{1}}\otimes|\circ\rangle\langle\circ|_{q_{5}}+X_{q_{1}}\otimes|\times\rangle\langle\times|_{q_{5}}\big{)}$
$=II\otimes|\circ\!\circ\rangle\langle\circ\!\circ|+IX\otimes|\circ\!\times\rangle\langle\circ\!\times|+XI\otimes|\times\!\circ\rangle\langle\times\!\circ|+XX\otimes|\times\!\times\rangle\langle\times\!\times|,$ (D.13)
where $q_{4}$ and $q_{5}$ are control qubits and $q_{0}$ and $q_{1}$ are target qubits. The action of $U_{2}$ on $I_{q_{0}}\otimes I_{q_{1}}\otimes|\xi\rangle_{2}$ gives
$U_{2}|\xi\rangle_{2}=\dfrac{1}{\sqrt{2}}\Big{(}IX\otimes|\circ\!\times\rangle+XI\otimes|\times\!\circ\rangle\Big{)},$ (D.14)
where, again, we omit $I_{q_{0}}\otimes I_{q_{1}}$ for short. After the interaction $U_{2}$, we apply the Hadamard gates ${\rm H}_{q_{4}}{\rm H}_{q_{5}}$ onto qubits $q_{4}$ and $q_{5}$ of the meter M2
${\rm H}_{q_{4}}{\rm H}_{q_{5}}U_{2}|\xi\rangle_{2}$ $=\dfrac{1}{2\sqrt{2}}\Big{(}IX\otimes\big{[}|\circ\rangle+|\times\rangle\big{)}\big{(}|\circ\rangle-|\times\rangle\big{)}+XI\otimes\big{(}|\circ\rangle-|\times\rangle\big{)}\big{(}|\circ\rangle+|\times\rangle\big{)}\Big{]}$
$=\dfrac{1}{2\sqrt{2}}\Big{[}\big{(}IX+XI\big{)}\otimes\big{(}|\circ\!\circ\rangle-|\times\!\times\rangle\big{)}+\big{(}IX-XI\big{)}\otimes\big{(}-|\circ\!\times\rangle+|\times\!\circ\rangle\big{)}\Big{]},$ (D.15)
where we used ${\rm H}|\circ\rangle=(|\circ\rangle+|\times\rangle)/\sqrt{2}$ and ${\rm H}|\times\rangle=(|\circ\rangle-|\times\rangle)/\sqrt{2}$. We next calculate the Kraus operators $N_{\nu}=\langle\nu|{\rm H}_{q_{4}}{\rm H}_{q_{5}}U_{2}|\xi\rangle_{2}$, where $\nu\in\{\circ\circ,\circ\times,\times\circ,\times\times\}$. We have
$N_{\circ\circ}$ $=-N_{\times\times}=\dfrac{1}{2\sqrt{2}}\big{(}IX+XI\big{)},$ (D.16)
$-N_{\circ\times}$ $=N_{\times\circ}=\dfrac{1}{2\sqrt{2}}\big{(}IX-XI\big{)},$ (D.17)
and the corresponding POVM yields
$E_{\circ\circ}$ $=E_{\times\times}=\dfrac{1}{4}\big{(}II+XX\big{)},$ (D.18)
$E_{\circ\times}$ $=E_{\times\circ}=\dfrac{1}{4}\big{(}II-XX\big{)},$ (D.19)
which satisfies $\sum_{\nu}E_{\nu}=II$. The probabilities yield
$P_{\circ\circ}$ $=P_{\times\!\times}=\langle\psi|E_{\circ\circ}|\psi\rangle=\dfrac{1}{4}\big{(}1+\cos\phi\sin 2\theta\big{)},$ (D.20)
$P_{\circ\times}$ $=P_{\times\circ}=\langle\psi|E_{\circ\times}|\psi\rangle=\dfrac{1}{4}\big{(}1-\cos\phi\sin 2\theta\big{)}.$ (D.21)
Finally, we get the expectation value $\langle XX\rangle$
$\langle XX\rangle=P_{\circ\circ}-P_{\circ\times}-P_{\times\circ}+P_{\times\!\times}=\cos\phi\sin 2\theta,$ (D.22)
as shown in Eq. (23) in the main text.
Appendix E Post-measurement state
In this section, we derive the final system state after measuring M1 and M2. The system state after the first measurement gives $\rho_{1}=\sum_{\mu}M_{\mu}|\psi\rangle\langle\psi|M_{\mu}^{\dagger}$. We first derive
$M_{\uparrow\uparrow}|\psi\rangle$ $=\dfrac{1}{\sqrt{2}}\big{(}|00\rangle\langle 00|+|11\rangle\langle 11|\big{)}\big{(}\cos\theta|00\rangle+e^{i\phi}\sin\theta|11\rangle\big{)}$
$=\dfrac{1}{\sqrt{2}}|\psi\rangle,$ (E.1)
and thus $M_{\uparrow\uparrow}|\psi\rangle\langle\psi|M^{\dagger}_{\uparrow\uparrow}=|\psi\rangle/2$. Similarly, we have $M_{\downarrow\downarrow}|\psi\rangle\langle\psi|M^{\dagger}_{\downarrow\downarrow}=|\psi\rangle/2$, and $M_{\uparrow\downarrow}|\psi\rangle\langle\psi|M^{\dagger}_{\uparrow\downarrow}=M_{\downarrow\uparrow}|\psi\rangle\langle\psi|M^{\dagger}_{\downarrow\uparrow}=0.$ Finally, we get
$\rho_{1}=\sum_{\mu}M_{\mu}|\psi\rangle\langle\psi|M_{\mu}^{\dagger}=|\psi\rangle\langle\psi|,$ (E.2)
Next, we derive the system state after the second measurement. It gives $\rho_{2}=\sum_{\nu}N_{\nu}\rho_{1}N_{\nu}^{\dagger}$. We first derive
$N_{\circ\circ}|\psi\rangle$ $=\dfrac{1}{2\sqrt{2}}\big{(}IX+XI\big{)}\big{(}\cos\theta|00\rangle+e^{i\phi}\sin\theta|11\rangle\big{)}$
$=\dfrac{1}{2\sqrt{2}}\big{(}\cos\theta+e^{i\phi}\sin\theta\big{)}\big{(}|01\rangle+|10\rangle\big{)},$ (E.3)
and thus
$\rho_{2}=\sum_{\nu}N_{\nu}\rho_{1}N_{\nu}^{\dagger}=\dfrac{1}{2}\begin{pmatrix}0&0&0&0\\ 0&1&\cos\phi\sin 2\theta&0\\ 0&\cos\phi\sin 2\theta&1&0\\ 0&0&0&0\end{pmatrix}.$ (E.4)
Appendix F Mixed state case
In this section, we provide detailed calculations for mixed-state cases. We need to derive Eqs.(32-35). We first recast the quantum state in matrix form as
$\rho=\begin{pmatrix}\frac{1-p}{4}&0&0&0\\ 0&\frac{1+p}{4}&-\frac{p}{2}&0\\ 0&-\frac{p}{2}&\frac{1+p}{4}&0\\ 0&0&0&\frac{1-p}{4}\end{pmatrix}.$ (F.1)
The probabilities give
$P_{\uparrow\uparrow}$ $={\rm Tr}\big{[}E_{\uparrow\uparrow}\rho\big{]}$
$=\dfrac{1}{2}{\rm Tr}\big{[}\big{(}|00\rangle\langle 00|+|11\rangle\langle 11\big{)}\rho\big{]}$
$=\dfrac{1}{2}\big{[}\langle 00|\rho|00\rangle+\langle 11|\rho|11\rangle\big{]}$
$=\dfrac{1-p}{4}.$ (F.2)
Similarly, we have $P_{\downarrow\downarrow}=\dfrac{1-p}{4}$, and
$P_{\uparrow\downarrow}=P_{\downarrow\uparrow}$ $=\dfrac{1}{2}(\langle 01|\rho|01\rangle+\langle 10|\rho|10\rangle)$
$=\dfrac{1+p}{4}.$ (F.3)
Finally, the expectation value for the measurement of $\langle ZZ\rangle$ gives
$\langle ZZ\rangle$ $=P_{\uparrow\uparrow}-P_{\uparrow\downarrow}-P_{\downarrow\uparrow}+P_{\downarrow\downarrow}$
$=\dfrac{1-p}{4}-\dfrac{1+p}{4}-\dfrac{1+p}{4}+\dfrac{1-p}{4}$
$=-p.$ (F.4)
[…]
$P_{\circ\circ}=P_{\times\times}$ $={\rm Tr}\big{[}E_{\circ\circ}\rho\big{]}$
$=\dfrac{1}{4}{\rm Tr}\big{[}(II+XX)\rho\big{]}$
$=\dfrac{1-p}{4},$ (F.5)
$P_{\circ\times}=P_{\times\circ}$ $={\rm Tr}\big{[}E_{\circ\times}\rho\big{]}$
$=\dfrac{1}{4}{\rm Tr}\big{[}(II+XX)\rho\big{]}$
$=\dfrac{1+p}{4}$ (F.6)
Then, the expectation value
$\langle XX\rangle=-p.$ (F.7)
The post-measurement states are given through $\rho_{1}=\sum_{\mu}M_{\mu}\rho M_{\mu}^{\dagger}$ and $\rho_{2}=\sum_{\nu}N_{\nu}\rho_{1}N_{\nu}^{\dagger}$ after measuring of M1 and M2, respectively. We first derive $M_{\mu}\rho M_{\mu}^{\dagger}$ for $\mu=\{\uparrow\uparrow,\uparrow\downarrow,\downarrow\uparrow,\downarrow\downarrow\}$, where
$M_{\uparrow\uparrow}\rho M_{\uparrow\uparrow}^{\dagger}=\dfrac{1}{2}\begin{pmatrix}\frac{1-p}{4}&0&0&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&0&0&\frac{1-p}{4}\end{pmatrix},$ (F.8)
and similar for $M_{\downarrow\downarrow}\rho M_{\downarrow\downarrow}^{\dagger}$, and
$M_{\uparrow\downarrow}\rho M_{\uparrow\downarrow}^{\dagger}=\dfrac{1}{2}\begin{pmatrix}0&0&0&0\\ 0&\frac{1+p}{4}&-\frac{p}{2}&0\\ 0&-\frac{p}{2}&\frac{1+p}{4}&0\\ 0&0&0&0\end{pmatrix},$ (F.9)
and similar for $M_{\downarrow\uparrow}\rho M_{\downarrow\uparrow}^{\dagger}$. Finally, we have
$\rho_{1}=\sum_{\mu}M_{\mu}\rho M_{\mu}^{\dagger}=\rho.$ (F.10)
Similar for $\rho_{2}$, we obtain
$\rho_{2}=\sum_{\nu}N_{\nu}\rho_{1}N_{\nu}^{\dagger}=\begin{pmatrix}\frac{1+p}{4}&0&0&-\frac{p}{2}\\ 0&\frac{1-p}{4}&0&0\\ 0&0&\frac{1-p}{4}&0\\ -\frac{p}{2}&0&0&\frac{1+p}{4}\end{pmatrix}.$ (F.11)